# Taming Data Challenges in ML-based Security Tasks: Lessons from Integrating Generative AI

### v1

@@ V-C Traditional data augmentation tools
We select two off-the-shelf non-GenAI baseline methods for comparison. These “traditional” approaches use rule-based techniques to weight existing samples and augment data under specific constraints. Implementation details are provided in Appendix Section -C.
@@
SMOTE [18].   SMOTE is a popular sample-conditioned technique that leverages linear interpolation between samples. Specifically, given an existing data sample from the real dataset, it generates synthetic data by sampling along the lines that connect the sample to its $k$-nearest neighbors. SMOTE does not scale well to high-dimensional imbalanced datasets [14]. A plethora of methods build on SMOTE [35]. DeepSMOTE [26] uses a latent-space interpolation via an encoder-decoder, but supports only image data, making it unsuitable for tabular tasks (Appendix -D). Despite its limitations, we include SMOTE for comparison, as it remains the most widely used data augmentation method [35].
@@
MC-CCR [58].   MC-CCR is a sample-conditioned technique with specific adaptations for the multi-class setting. Before generating synthetic data, MC-CCR performs a “cleaning” stage, where samples with “wrong” labels are repositioned by moving samples from other classes which are close to samples of the target class. Re-positioning is achieved by translating the mislabeled samples to the surface of a sphere around it. If many wrong labeled samples are re-positioned around a data sample, it is considered “difficult”. MC-CCR thus generates synthetic data by sampling within the sphere, with more synthetic data being generated for “difficult” samples. Additional details are in Appendix Section -E.
@@
Other methods.    We do not evaluate other SMOTE-based oversampling methods, as MC-CCR has been shown to outperform both other multi-class oversampling techniques and simple GAN-based generators in highly imbalanced settings [26]. We also exclude MC-RBO [59], a comparable method, since unlike MC-CCR, it lacks a cleaning strategy to remove outliers before augmentation in multi-class scenarios.
V-D Nimai: Enabling controlled data synthesis
@@
Recall that existing GenAI schemes are class-conditional synthesis schemes. To alleviate complex biases within a class, the defender may require more control over the data generation process. Class-conditional synthesis or uncontrolled synthesis that creates many synthetic samples to broadly mimic real samples may not be the best strategy in all cases. We propose a new GenAI scheme called Nimai that provides highly controlled data synthesis through a sample-conditioning scheme. Such an approach would help target regions of the data manifold that are underrepresented. That said, Nimai also supports class-conditional synthesis. In addition, Nimai provides automated mechanisms to efficiently tune its hyper-parameters. Recall that among the existing GenAI schemes, only TabDDPM offers strategies for hyper-parameter tuning when fitting the model to diverse datasets. We emphasize that the goal of Nimai is not to provide a “one-fits-all” solution that supersedes the GenAI schemes discussed in Section V-B; rather, we aim to understand the impact of using sample-conditioning, compared to a class-conditional approach, for security tasks.
Architecture and training.    Nimai is a VAE-based GenAI model. Unlike GANs, VAEs are easier to train due to their tractable likelihood loss [17], and their likelihood maximization encourages greater sample diversity [100]. Compared to Diffusion models, VAEs offer faster sampling, avoiding the costly iterative denoising process.
We aim to create synthetic samples in the “neighborhood” of existing samples (e.g., in minority classes), thus improving generalization performance. Our key idea is to use a discrete latent space for the VAE model to controllably produce synthetic samples that generalize better. In comparison, existing VAE-based methods like TVAE and TabSyn encode features into a continuous latent space, which can result in prior holes and posterior collapse [7, 68] impacting performance.
Recently, discrete latent codes have driven success across tasks, especially in text generation [53], as well as image and video generation [9]. For instance, LLMs model distributions over discrete variables, enabling highly controlled text generation [53, 28].
Nimai adapts a Vector Quantized Variational AutoEncoder (VQ-VAE) [101] to enable a discrete latent space. We convert the 2D discrete latent tokens used for images into 1D latent vectors suitable for tabular data.
Figure 3: Architecture of Nimai framework.
@@
Figure 3 depicts Nimai’s architecture. We implement a discrete latent space by maintaining a shared embedding space or a codebook that contains as many embedding vectors as the size of the discrete latent space (e.g., 1 of $K$ categorical values). Transformer blocks are used as encoder/decoder. Tabular data is normalized between 0 and 1 to support stable encoder/decoder learning. An input tabular sample, alongwith its class label, is first converted to a continuous representation by the encoder block. This continuous output is discretized by identifying the nearest embedding vectors in the (trainable) codebook, i.e., discretized by using the indices of the nearest vectors in the codebook. The decoder receives this discrete latent vector as input. This model is trained to bound the log-likelihood using the evidence lower bound (ELBO), in addition to optimizing other loss terms that maintain the codebook vectors closer to the encoder’s output, and ensure the encoder commits to an embedding. Nimai’s training objectives include the following loss terms:
Reconstruction loss ($\mathcal{L}_{\text{recon}}$) This term ensures accurate reconstruction in feature space using mean squared error (MSE). For input $x$ and reconstruction $\hat{x}$, $\mathcal{L}_{\text{recon}}$ is defined in Equation 1.
$\mathcal{L}_{\text{recon}}=\|x-\hat{x}\|^{2}$ (1)
Commitment Loss ($\mathcal{L}_{\text{commit}}$) This loss aligns encoder outputs with nearest codebook embeddings. As shown in Equation 2, $z_{e}(x)$ is the encoder output, $e$ the closest embedding, $sg(\cdot)$ stops gradients, and $\beta$ scales the loss.
$\mathcal{L}_{\text{commit}}=\beta\|z_{e}(x)-\text{sg}(e)\|^{2}$ (2)
Embedding Loss ($\mathcal{L}_{\text{embed}}$) This term adjusts embedding vectors to match encoder outputs. Equation 3 shows the embedding loss, where the stop-gradient $sg(\cdot)$ prevents gradients from flowing to the encoder output.
$\mathcal{L}_{\text{embed}}=\|\text{sg}(z_{e}(x))-e\|^{2}$ (3)
Nimai’s objective, shown in Equation 4, combines all three losses, weighted by $\alpha$. Early stopping on validation loss prevents overfitting.
$\mathcal{L}=\mathcal{L}_{\text{recon}}+\alpha\left(\mathcal{L}_{\text{embed}}+\mathcal{L}_{\text{commit}}\right)$ (4)
@@
We develop Masked Token Modeling (MTM) to build the prior. This is inspired by work in the image [17] and natural language domain [52]. More concretely, to perform sample-conditioning, we use the encoder to project a sample in the input feature space to the latent space, and then mask, i.e., “blank-out” a subset of the latent features. Our MTM technique takes a feature vector and its class label as input, and predicts masked feature values using surrounding context. It is implemented using the BERT model [52]. We fill-in (infer) new features in a non-autoregressive manner in a constant number of timesteps. MTM’s objective in Equation 5 uses cross-entropy loss to estimate the log-likelihood of quantized latent tokens. Here, $\mathbf{y}_{i}$ is a latent feature, $\mathbf{Y}_{\text{mask}}$ a masked token filled by MTM, and $\mathbf{c}$ the class label.
$\mathcal{L}_{\text{MTM}}=\mathbb{E}_{(\mathbf{y},c)\sim D}\left[-\sum_{i}\log p\left(\mathbf{y}_{i}\mid\mathbf{Y}_{\text{mask}},c\right)\right]$ (5)
At generation time (Stage III of Figure 3), a latent feature vector with the “blanks-filled-in” (synthesized vector) with its class label is passed through the decoder, mapping it to a nearby sample in the input feature space.
@@
Note that sample-conditioning via MTM reduces to the class-conditioned case when all the latent features are masked out. As such, Nimai inherently allows for interpolating between sample-conditioned or class-conditioned sampling, i.e., treats the latent feature masking ratio as a hyper-parameter that can be chosen using the validation set. We refer to the sample-conditioned and class-conditioned versions of Nimai as Nimai-S and Nimai-C, respectively. Nimai uses a hyperparameter search with the ASHA algorithm[65] across various architectural choices (Table IX, Appendix), making it more adaptable to diverse security data. Further details are provided in Appendix Section -F.
VI Evaluation Setup
@@
Feature Pre-processing.    Tabular data in some of our tasks contains mixed data types like continuous, and discrete-valued features. Training generative models on mixed feature types is a non-trivial task [116]. We leave this aspect for future work. This work focuses on alleviating the data challenges mentioned in Section II. In our evaluation, we standardize all feature values between 0 and 1, irrespective of the data type. We find that this scaling has minimal impact on model performance in 6/7 datasets where the end defense classifier is a tree-based model e.g., XGBoost, LightGBM, etc. The exception was IoT, where a non-tree based classifier, i.e., Logistic Regression is used, the performance increases from 46% to 74.16% macro F-score after data standardization. Note that despite the improvement, all augmentation schemes (both GenAI and traditional approaches) are evaluated on the same standardized dataset, allowing relative performance comparison.
@@
Performance metrics.    For each task in Section IV, we use the same training configuration (including the proposed classifier and associated hyperparameters) with and without data augmentation. We measure the macro average F-score for all tasks, except for Tor and BGP, where we report the F-score for the malicious class (see Section IV). For any given $N-$class classifier, let $P_{i}$ and $R_{i}$ represent the Precision and Recall on the $i^{\text{th}}$ class. The macro average F-score is then given by:
$F=\dfrac{2}{N}\sum_{i=1}^{N}\frac{P_{i}\cdot R_{i}}{P_{i}+R_{i}}$ (6)
@@
which may be viewed as an unweighted average of individual F-scores across each of the $N$ classes. We focus on evaluating relative gains in classifier performance. Gain may also be negative if there is performance degradation after data augmentation. Let $F_{real}$ and $F_{aug}$ represent the F-scores of the classifier before and after data augmentation. We then measure relative gain $\Delta G$ as:
$\Delta G=\frac{F_{aug}-F_{real}}{F_{real}}$ (7)
@@
When measuring $\Delta G$, we repeat experiments across 10 trials with different random seeds and compute the mean and standard deviation of F-scores. Note that for two tasks BODMAS and MS-Malware, when computing $\Delta G$, we omit classes with too few test set samples ($\leq$ 20). Higher, positive values of $\Delta G$ indicate better performance from data augmentation. Note that our experimental setup allows us to directly attribute any performance gain (or degradation), i.e., $\Delta G$, to the synthetic data used for the augmentation. Our setup also enables comparison across data augmentation schemes (GenAI and non-GenAI schemes) for a given task. For the same task, all data augmentation schemes share the same classifier training setup, except for the differences in the augmented datasets.
@@
Identifying unreliable measurements.    We use the Coefficient of Variation (CV) [3] metric to identify unreliable measurements of $\Delta G$. CV is computed as the ratio of the absolute standard deviation to the absolute mean. Notably, we observe cases where the CV for $\Delta G$ exceeds 1 — in such cases, we consider the data augmentation scheme to have a high performance dispersion, and deem the measurement to be unreliable. We focus our discussions primarily on reliable schemes and label unreliable schemes as such.
@@
Statistical tests.    In certain cases, comparing reliable data augmentation schemes using a simple ordering of their mean $\Delta G$ scores can be problematic, e.g.,, when a scheme with larger mean is accompanied by a larger standard deviation. To address these ambiguities, we compare such $\Delta G$ rankings using statistical tests. We use the non-parametric Kruskal-Wallis [60] test to determine if there is indeed a ranking of the schemes. We chose this test because it does not require one to satisfy the assumption of homogeneity of variance [72]. When Kruskal-Wallis test ranks differ, we apply Dunn’s [33] test for post-hoc pairwise ranking comparisons.
VII Evaluation: Pre-Deployment
@@
The focus here is on improving performance via data augmentation before deploying the classifier (see Figure 1). Evaluation using the 7 data-challenged tasks (Section II) are divided into two parts: (1) cases where GenAI schemes demonstrate clear potential, and (2) cases where most GenAI schemes fall short, i.e., failed to significantly boost performance. Our analysis sheds light on the strengths and limitations of GenAI schemes. We list the quantity of synthetic samples added per task in Appendix Table XII and the training times for all GenAI schemes in Appendix Table VII.
@@
TVAE  2.02 (0.66) $\uparrow$ -20.83 (3.22) $\downarrow$ 11.32 (5.86) $\uparrow$
@@
TABLE II: Mean $\Delta G$ with standard deviation for IoT, BGP, and MS-Malware tasks, where GenAI shows potential. $\uparrow$ and $\downarrow$ indicate positive and negative gains, respectively. ‘-’ indicates unreliable gains (CV $\geq$ 1), and ‘$\times$’ indicates model failures.
@@
TABLE III: Mean $\Delta G$ with standard deviation for BODMAS task for 2, 3, 5, 6, 7, 9, 10 & 11 months. $\uparrow$ and $\downarrow$ indicate positive and negative gains, respectively. ‘-’ indicates unreliable gains (CV $\geq$ 1). GReaT and REaLTabFormer’s failures are not shown here. Results for months 1, 4, 8 & 12 where we did not observe any performance gains are in Table VI (Appendix).
VII-A Cases where GenAI shows potential
@@
IoT.    The IoT task suffers from limited or biased data challenges (a)-(c) (Section II), with a notable challenge of high class imbalance ratio of 38:1 (Section IV). Recall that the task has a total of 43,866 training samples of which only 1,116 are attack samples. Results are in Table II.
Finding 1.   Sample-conditioned GenAI schemes have the potential to boost performance in highly imbalanced class settings in binary classification tasks, outperforming traditional data augmentation schemes.
We observe that the best-performing approaches use sample-conditioning. Nimai-S achieves a mean $\Delta G$ of 17.37% outperforming all the schemes. Class-conditional GenAI approaches also demonstrate positive $\Delta G$, notably TVAE, CTAB-GAN+, and Nimai-C. One possible reason for CTAB-GAN+’s improved performance over Nimai-C and TVAE is due to its special loss terms designed to handle imbalanced data. We also note that Nimai-C gives a mean $\Delta G$ of 3.93%, which is greater than that of TVAE (mean $\Delta G$ of 2.02%). This could be due to the quantized space used by Nimai. The traditional scheme SMOTE is able to beat most GenAI schemes, but not Nimai-S. This again highlights the potential for GenAI to further push the boundaries of data augmentation. TabSyn and MC-CCR exhibit unreliable behavior as per our definition in Section VI.
Failure cases. Only 3/6 existing GenAI tools (i.e., excluding Nimai) were successfully instantiated on the IoT task data. We find that TabDDPM, GReaT and REaLTabFormer are unsuccessful at producing usable classifiers. Specifically, GReaT produces an inordinately large natural language representation of the 273 IoT features, requiring about 5,714 hours to train for one epoch using an NVIDIA A100 GPU (See Appendix Table VII). REaLTabFormer fails to train on IoT as it does not reach stopping criteria even after 40 hours of training. TabDDPM faces different problems and suffers from mode collapse, i.e., it is unable to produce samples from the minority class. Training on such synthetic samples (i.e., only from the majority class) yields an unusable classifier that labels all inputs as the majority class.
BGP.    The BGP task suffers from limited or biased data challenges (a)-(c) (Section II), with the notable challenge being the extremely limited training data. The training data has only 180 samples across 2 classes, with only 17 samples in the malicious class. Results are in Table II.
@@
Four GenAI schemes, TVAE, TabSyn, GReaT and REaLTabFormer produce performance degradation. These schemes are not best suited for extremely data-challenged settings. Despite using a large pre-trained LLM backend (i.e., Distill GPT-2 [83]), GReaT and REaLTabFormer fail to demonstrate any gains, again highlighting that LLM-based schemes require further engineering advances before they can be practically adopted. The traditional scheme, MC-CCR also results in a performance degradation. We posit that, in high-class imbalance cases, the minority point spheres become extremely small, leading to points highly similar to the original data samples, thereby likely leading to overfitting.
@@
BODMAS.   The BODMAS task has limited and bias data challenges (a), (b) and (d), along with the notable challenge of concept drift (Section II). Concept drift can be observed by the drop in macro F-score across the 12 test months, especially for months 5 and 6 (see Figure 2). Data from month-0 is used to train a GDBT classifier and testing is done on the subsequent 12 months of malware traces. The results are shown in Table III and Table VI (Appendix).
@@
Mitigating concept drift is a challenging problem [6]. We find that practitioners can significantly improve the performance of security classifiers before deployment by integrating GenAI. For 8/12 months, we observe a positive gain in performance, when using a GenAI scheme. In fact, we achieve over 59% mean $\Delta G$ for month 5, where the largest impact of the concept drift was observed (Figure 2). This is a substantial result. The results are shown in Table III. For the remaining 4 months, none of the data augmentation schemes demonstrate a gain in performance (see Table VI in the Appendix). This is expected, given the challenging nature of the problem. However, achieving performance gains in 8/12 mounts is a notable result. For the remainder of this section, we analyze the 8 months shown in Table III.
@@
Finding 3.   Class-conditioned GenAI approaches are well suited for tasks where there is concept drift. A key finding is that class-conditioned GenAI approaches are a better fit for tasks with concept drift, compared to sample-conditioned schemes. Class-conditioning can better explore novel regions of the data manifold, where distribution shift may have occurred. This is because we are not explicitly focusing on generating samples near the vicinity of existing samples. Our class-conditioned Nimai-C achieves the best performance in 6 out of the 8 months (Table III), i.e., months 2, 3, 5, 9, 10, and 11. Detailed statistical test results to arrive at this conclusion are discussed in the Appendix Section -G. On the other hand, our sample-conditioned scheme Nimai-S produces unreliable performance estimates in 5 out of the 8 months, and trails behind other GenAI schemes in 2 out of the remaining 3 months. Other notable schemes are TVAE and TabDDPM, which perform similar to Nimai-C (no statistically significant difference) in months 9 and 10, respectively. TVAE outperforms all schemes only in month 6.

### v2

@@ 5.3. Traditional data augmentation tools
We select two off-the-shelf non-GenAI baseline methods for comparison. These “traditional” approaches use rule-based techniques to weight existing samples and augment data under specific constraints. Implementation details are provided in Section C.
@@
SMOTE (smote).   SMOTE is a popular sample-conditioned technique that leverages linear interpolation between samples. Specifically, given an existing data sample from the real dataset, it generates synthetic data by sampling along the lines that connect the sample to its $k$-nearest neighbors. SMOTE does not scale well to high-dimensional imbalanced datasets (blagus2013smote). A plethora of methods build on SMOTE (fernandez2018smote). One of them, DeepSMOTE (dablain2022deepsmote) uses a latent-space interpolation via an encoder-decoder, but supports only image data, making it unsuitable for tabular tasks (Appendix D). Despite its limitations, we include SMOTE for comparison, as it remains the most widely used data augmentation method (fernandez2018smote).
@@
MC-CCR (mc-ccr).   MC-CCR is a sample-conditioned technique with specific adaptations for the multi-class setting. Before generating synthetic data, MC-CCR performs a “cleaning” stage, where samples with “wrong” labels are repositioned by moving samples from other classes which are close to samples of the target class. Re-positioning is achieved by translating the mislabeled samples to the surface of a sphere around it. If many wrong labeled samples are re-positioned around a data sample, it is considered “difficult”. MC-CCR thus generates synthetic data by sampling within the sphere, with more synthetic data being generated for “difficult” samples. Additional details are in Appendix E.
@@
Other methods.    We do not evaluate other SMOTE-based oversampling methods, as MC-CCR has been shown to outperform both other multi-class oversampling techniques and simple GAN-based generators in highly imbalanced settings (dablain2022deepsmote). We also exclude MC-RBO (krawczyk2019radial), since unlike MC-CCR, it lacks a cleaning strategy to remove outliers before augmentation in multi-class scenarios.
Figure 3. Architecture of Nimai framework.:
5.4. Nimai: Enabling controlled data synthesis
@@
Recall that existing GenAI schemes are class-conditional synthesis schemes. To alleviate complex biases within a class, the defender may require more control over the data generation process. Class-conditional synthesis or uncontrolled synthesis that creates many synthetic samples to broadly mimic real samples may not be the best strategy in all cases. We propose a new GenAI scheme called Nimai that provides highly controlled data synthesis through a sample-conditioning scheme. Such an approach would help target regions of the data manifold that are underrepresented. That said, Nimai also supports class-conditional synthesis. In addition, Nimai provides automated mechanisms to efficiently tune its hyper-parameters. Recall that among the existing GenAI schemes, only TabDDPM offers strategies for hyper-parameter tuning when fitting the model to diverse datasets. We emphasize that the goal of Nimai is not to provide a “one-fits-all” solution that supersedes the GenAI schemes discussed in Section 5.2; rather, we aim to understand the impact of using sample-conditioning, compared to a class-conditional approach, for security tasks.
Architecture and training.    Nimai is a VAE-based GenAI model. Nimai adapts a Vector Quantized Variational AutoEncoder (VQ-VAE) (vqvae) to enable a discrete latent space. We convert the 2D discrete latent tokens used for images into 1D latent vectors suitable for tabular data. A key innovation lies in integrating a BERT-based masked-token-model (MTM) in the latent space, enabling sample-based generation with controllable variation via a masking ratio and supporting class-conditional generation. Unlike GANs, VAEs are easier to train due to their tractable likelihood loss (chang2022maskgit), and their likelihood maximization encourages greater sample diversity (van2016conditional). Compared to Diffusion models, VAEs offer faster sampling, avoiding the costly iterative denoising process. We aim to create synthetic samples in the “neighborhood” of existing samples (e.g., in minority classes), thus improving generalization performance. Our key idea is to use a discrete latent space for the VAE model to controllably produce synthetic samples that generalize better. In comparison, existing VAE-based methods like TVAE and TabSyn encode features into a continuous latent space, which can result in prior holes and posterior collapse (aneja2021contrastive; lucas2019understanding) impacting performance. Recently, discrete latent codes have driven success across tasks, especially in text generation (keskar2019ctrl), as well as image and video generation (bai2024sequential). For instance, LLMs model distributions over discrete variables, enabling highly controlled text generation (keskar2019ctrl; dathathri2019plug).
@@
Figure 3 depicts Nimai’s architecture. We implement a discrete latent space via a shared embedding codebook with $K$ vectors (one per categorical value). Transformer blocks serve as encoder/decoder, and tabular data is normalized to [0,1] for stable learning. An input tabular sample, along with its class label, is first converted to a continuous representation by the encoder block. This continuous output is discretized by identifying the nearest embedding vectors in the (trainable) codebook, i.e., discretized by using the indices of the nearest vectors in the codebook. The decoder receives this discrete latent vector as input. The model is trained to bound log-likelihood via the evidence lower bound (ELBO), while also optimizing losses that align codebook vectors with encoder outputs and enforce encoder commitment. A full description of the training objective is provided in Appendix I.
@@
We develop Masked Token Modeling (MTM) to build the prior. This is inspired by work in the image (chang2022maskgit) and natural language domain (kenton2019bert). More concretely, to perform sample-conditioning, we use the encoder to project a sample in the input feature space to the latent space, and then mask, i.e., “blank-out” a subset of the latent features. Our MTM technique takes a feature vector and its class label as input, and predicts masked feature values using surrounding context. It is implemented using the BERT model (kenton2019bert). We fill-in (infer) new features in a non-autoregressive manner in a constant number of timesteps. A detailed description of the MTM objective is provided in Appendix J. At generation time (Stage III of Figure 3), a latent feature vector with the “blanks-filled-in” (synthesized vector) with its class label is passed through the decoder, mapping it to a nearby sample in the input feature space. We ensure that synthetic data preserve the value ranges of the original feature space.
@@
Note that sample-conditioning via MTM reduces to the class- conditioned case when all the latent features are masked out. As such, Nimai inherently allows for interpolating between sample-conditioned or class- conditioned sampling, i.e., treats the latent feature masking ratio as a hyper-parameter that can be chosen using the validation set. We refer to the sample-conditioned and class-conditioned versions of Nimai as Nimai-S and Nimai-C, respectively. Nimai uses a hyperparameter search with the ASHA algorithm(li2020system) across various architectural choices (Table 10, Appendix), making it adaptable to diverse security data. Details are provided in Appendix F.
6. Evaluation Setup
@@
Feature Pre-processing.    Tabular data in some of our tasks contains mixed data types like continuous and discrete-valued features. Training generative models on mixed feature types is a non-trivial task (ctabgan). We leave this aspect for future work. This work focuses on alleviating the data challenges mentioned in Section 2. In our evaluation, we standardize all feature values between 0 and 1, irrespective of the data type. All features are normalized to [0,1] with a MinMax scaler (using clipping to keep held-out values in range). Normalized real data trains GenAI schemes, and the same scaler ensures that all synthetic data also lies within [0,1]. We find that this scaling has minimal impact on model performance in 6/7 datasets where the end defense classifier is a tree-based model e.g., XGBoost, LightGBM. The exception was IoT, where a non-tree based classifier, i.e., Logistic Regression is used, the performance increases from 46% to 74.16% macro F-score after data standardization. Note that despite the improvement, all augmentation schemes (GenAI and traditional approaches) are evaluated on the same standardized dataset, i.e., relative performance comparison.
@@
Performance metrics.    For each task in Section 4, we apply the same training setup (classifier and hyperparameters) with and without augmentation. We measure the macro average F-score for all tasks, except for Tor and BGP, where we report the F-score for the malicious class (see Section 4). For any given NN-class classifier, let $P_{i}$ and $R_{i}$ represent the Precision and Recall on the $i^{\text{th}}$ class. The macro average F-score is then given by:
(1) $F=\dfrac{2}{N}\sum_{i=1}^{N}\frac{P_{i}\cdot R_{i}}{P_{i}+R_{i}}$
@@
which may be viewed as an unweighted average of individual F-scores across each of the $N$ classes. We focus on evaluating relative gains in classifier performance. Gain may also be negative if there is performance degradation after data augmentation. Let $F_{real}$ and $F_{aug}$ represent the F-scores of the classifier before and after data augmentation, respectively. We then measure relative gain $\Delta G$ as:
(2) $\Delta G=\frac{F_{aug}-F_{real}}{F_{real}}$
@@
When measuring $\Delta G$, we repeat experiments across 10 trials with different random seeds and compute the mean and standard deviation of F-scores. Note that for two tasks BODMAS and MS-Malware, when computing $\Delta G$, we omit classes with too few test set samples ($\leq$ 20). Higher, positive values of $\Delta G$ indicate better performance from data augmentation. Our setup enables direct attribution of performance changes (i.e., $\Delta G$) to the synthetic data used for augmentation and supports fair comparison across GenAI and non-GenAI schemes. For each task, all augmentation methods use the same classifier configuration, differing only in the augmented datasets.
@@
Identifying unreliable measurements.    We use the Coefficient of Variation (CV) (abdi2010coefficient) to identify unreliable $\Delta G$ measurements, computed as the ratio of standard deviation to mean. When CV exceeds 1, we deem the augmentation scheme to have high dispersion and the measurement unreliable. We focus our discussions primarily on reliable schemes and label unreliable schemes as such.
@@
Statistical tests.    In certain cases, comparing reliable data augmentation schemes using a simple ordering of their mean $\Delta G$ scores can be problematic, e.g., when a scheme with larger mean is accompanied by a larger standard deviation. To resolve these ambiguities, we compare $\Delta G$ rankings using the non-parametric Kruskal–Wallis test (kruskal1952use), which does not assume homogeneity of variance (montgomery2019design). When Kruskal-Wallis test ranks differ, we apply Dunn’s (dunn1964multiple) test for post-hoc pairwise ranking comparisons.
7. Evaluation: Pre-Deployment
@@
The focus here is on improving performance via data augmentation before deploying the classifier (see Figure 1(a)). Evaluation using the 7 data-challenged tasks (Section 2) are divided into two parts: (1) cases where GenAI schemes demonstrate clear potential, and (2) cases where most GenAI schemes fall short, i.e., failed to significantly boost performance.
We list the quantity of synthetic samples added per task in Table 9 (Appendix) and the training times for all GenAI schemes in Table 8 (Appendix).
@@
TVAE  2.02 (0.66) $\uparrow$ -20.83 (3.22) $\downarrow$ 11.46 (5.88) $\uparrow$
@@
Table 2. Mean $\Delta G$ with standard deviation for IoT, BGP, and MS-Malware tasks, where GenAI shows potential. $\uparrow$ and $\downarrow$ indicate positive and negative gains, respectively. ‘-’ indicates unreliable gains (CV $\geq$ 1), and $\times$’ indicates model failures.
@@
Table 3. Mean $\Delta G$ with standard deviation for BODMAS task for 2, 3, 5, 6, 7, 9, 10 & 11 months. $\uparrow$ and $\downarrow$ indicate positive and negative gains, respectively. ‘-’ indicates unreliable gains (CV $\geq$ 1). GReaT and REaLTabFormer’s failures are not shown here. Results for months 1, 4, 8 & 12 are in Table 7 (Appendix) where no gains were observed.
7.1. Cases where GenAI shows potential
@@
IoT.    The IoT task suffers from limited or biased data challenges (a)-(c) (Section 2), with a notable challenge of high class imbalance ratio of 38:1 (Section 4). Recall the task has a total of 43,866 training samples of which only 1,116 are attack samples. Results are in Table 2.
Finding 1.   Sample-conditioned GenAI schemes have the potential to boost performance in highly imbalanced class settings in binary classification tasks, outperforming traditional data augmentation schemes. We observe that the best-performing approaches use sample-conditioning. Nimai-S achieves a mean $\Delta G$ of 17.37% outperforming all the schemes. Class-conditional GenAI approaches also demonstrate positive $\Delta G$, notably TVAE, CTAB-GAN+, and Nimai-C. One possible reason for CTAB-GAN+’s improved performance over Nimai-C and TVAE is due to its special loss terms designed to handle imbalanced data. Nimai-C achieves a mean $\Delta G$ of 3.93%, surpassing TVAE (2.02%), likely due to the quantized space of Nimai. While SMOTE outperforms most GenAI schemes, it falls short of Nimai-S, underscoring GenAI’s potential to advance augmentation. In contrast, TabSyn and MC-CCR show unreliable behavior.
Failure cases. Only 3/6 existing GenAI tools (i.e., excluding Nimai) were successfully instantiated on the IoT task data. We find that TabDDPM, GReaT, and REaLTabFormer fail to produce usable classifiers. In particular, GReaT generates an excessively large language representation of the 273 IoT features, requiring about 5,714 hours for a single epoch on an NVIDIA A100 GPU (See Table 8 (Appendix)). REaLTabFormer fails to train on IoT as it does not reach stopping criteria even after 40 hours of training. TabDDPM faces different problems and suffers from mode collapse, i.e., it is unable to produce samples from the minority class. Training on such synthetic samples (i.e., only from the majority class) yields an unusable classifier that labels all inputs as the majority class.
BGP.    The BGP task suffers from limited or biased data challenges (a)-(c) (Section 2), with the notable challenge being the extremely limited training data. Training data contains 180 samples across 2 classes, with only 17 samples in the malicious class. Results are in Table 2.
@@
Four GenAI schemes, TVAE, TabSyn, GReaT and REaLTabFormer produce performance degradation. These schemes are not best suited for extremely data-challenged settings. Despite using a large pre-trained LLM backend (i.e., Distill GPT-2 (sanh2019distilbert)), GReaT and REaLTabFormer fail to demonstrate any gains, again highlighting that LLM-based schemes require further engineering advances before they can be practically adopted. The traditional scheme, MC-CCR also results in a performance degradation. We posit that, in high-class imbalance cases, the minority point spheres become extremely small, leading to points highly similar to the original data samples, thereby likely leading to overfitting.
@@
BODMAS.   The BODMAS task has limited and bias data challenges (a), (b) and (d), along with the notable challenge of concept drift (Section 2). Concept drift can be observed by the drop in macro F-score across the 12 test months, especially for months 5 and 6 (see Figure 2). Data from month-0 is used to train a GDBT classifier and testing is done on the subsequent 12 months of malware traces. The results are shown in Table 3 and Table 7 (Appendix).
@@
Mitigating concept drift is a challenging problem (andresini2021insomnia). We find that practitioners can significantly improve the performance of security classifiers before deployment by integrating GenAI. For 8/12 months, we observe a positive gain in performance, when using a GenAI scheme. In fact, we achieve over 59% mean $\Delta G$ for month 5, where the largest impact of the concept drift was observed (Figure 2). This is a substantial result. The results are shown in Table 3. For the remaining 4 months, none of the data augmentation schemes demonstrate a gain in performance (see Table 7 in the Appendix). This is expected, given the challenging nature of the problem. However, achieving performance gains in 8/12 months is a notable result. For the remainder of this section, we analyze the 8 months shown in Table 3.
@@
Finding 3.   Class-conditioned GenAI approaches are well suited for tasks where there is concept drift. A key finding is that class-conditioned GenAI approaches are a better fit for tasks with concept drift, compared to sample-conditioned schemes. Class-conditioning can better explore novel regions of the data manifold, where distribution shift may have occurred. This is because we are not explicitly focusing on generating samples near the vicinity of existing samples. Our class-conditioned Nimai-C achieves the best performance in 6 out of the 8 months (Table 3), i.e., months 2, 3, 5, 9, 10, and 11. Detailed statistical test results to arrive at this conclusion are discussed in Section G.
On the other hand, our sample-conditioned scheme Nimai-S produces unreliable performance estimates in 5 out of the 8 months, and trails behind other GenAI schemes in 2 out of the remaining 3 months. Other notable schemes are TVAE and TabDDPM, which perform similar to Nimai-C (no statistically significant difference) in months 9 and 10, respectively. TVAE outperforms all schemes only in month 6. Traditional schemes (also sample-conditioned), SMOTE and MC-CCR perform poorly for this task. For 7 out of the 8 months, these schemes produce either negative gains or unreliable performance estimates. This highlights the limitations of traditional methods to mitigate concept drift.